A species of bacteria doubles every 2 hours, so that a culture of 200 bacteria would grow to 400 after 2 hours and 800 after 4 hours. If there were b bacteria initially, which of the following expressions correctly describes the amount of bacteria present after one day?
Answer by Boreal(15235)

A=Pe^kt
A/P=2=e^kt
ln2=kt, and t=2 hours
ln2=2k
k=(1/2)ln 2.
After one day, which is 24 hours,
A=b*e^(12*ln 2)
A=4096*b